OpenStudy (anonymous):

This is going to require a bit of time ;) Synthetically divide: f(x) = 2x^3 + 4x^2 + 2x - 3 – 4 divided by x + 2

OpenStudy (amistre64):

not much time really, its just adding and multiplying

OpenStudy (anonymous):

I just need it set up :/ I know you put 2 in the box, what else?

OpenStudy (amistre64):

3 2 1 0 2x^3 + 4x^2 + 2x - 3 – 4; this end is wierd, is it a typo?

OpenStudy (amistre64):

and my setup tends to be different than a textbook ..... a personal preference type thing

OpenStudy (anonymous):

Nope, it has to have a remainder of 4. SHould I change it to just -4 instaead of -3-4?

OpenStudy (amistre64):

well, -3-4 = -7 ... but are you saying the division needs to end in a -4?

OpenStudy (anonymous):

Yes, it has to have a remainder of 4

OpenStudy (anonymous):

Sorry, a remainder of -4

OpenStudy (amistre64):

first thing i do is make sure all the x parts are there .... high to low in sequential order 3 2 1 0 ; they seem to be accounted for 2x^3 + 4x^2 + 2x - 3 – 4 I dont remove the x parts, i use them as a visual cue, since they are redundant anyhow. Since we are dividing by x+2, then when x=-2 we get a zero, so lets divide this by the -2 2x^3 + 4x^2 + 2x - 3 – 4 0 <-- add this row -------------------------- -2 | <-- multiply this row 2x^3 + 4x^2 + 2x - 3 – 4 0 -4 0 <-- add this row -------------------------- -2 | 2 0 2 <-- multiply this row like that, ill let you work out the last parts

OpenStudy (anonymous):

Wow. I so got it now! That's not even sarcasm, I'll write this down as a reference later. Thanks for your help *medal*

OpenStudy (amistre64):

good luck :)

OpenStudy (amistre64):

just as a wrap up, the numbers in the bottom row are the coeffs of the "qoutient" polynomial; which is where my visual cue of the x parts helps me out. 2x^3 + 4x^2 + 2x - 3 – 4 0 -4 0 -------------------------- -2 | 2 0 2 (......) x^2 x c R \[\frac{2x^3 + 4x^2 + 2x - 3 – 4}{x+2}~=~ 2x^2 + 2 + \frac{R}{x+2}\]